OpenStudy (calculusxy):

The average weight of three dogs is 55 pounds, and the average weight of five cats is 9 pounds. What is the average weight, in pounds, of all eight animals?

OpenStudy (calculusxy):

@perl

OpenStudy (calculusxy):

@Nnesha

OpenStudy (anonymous):

Average weight of the animals: \(\sf \dfrac{(number\ of\ dogs\ \times avg\ weight\ of\ 3\ dogs)+(number\ of\ cats\ \times avg\ weight\ of\ 5\ cats)}{total\ number\ of\ animals}\)